# Robust Identity Perceptual Watermark Against Deepfake Face Swapping

### v2

@@ 3.3 Watermark Encoding
Watermark Features. As for the binary identity perceptual watermark $m$ of length $l$, we first increase the length linearly via a fully connected layer followed by dimension expansion via shape rearrangement and then leverage a convolutional block to elaborate the channel perspective. After that, we devise $N_{\textrm{enc}}$ consecutive diffusion blocks (denoted as Diff. Block) utilizing deconvolutions and smoothly diffuse the feature elements that represent $m$ to the same dimension as the width and height of $I$. Similarly, a series of $N_{\textrm{enc}}$ SE Blocks is promoted for the preparation of watermarking.
Image Reconstruction. Lastly, features of $I$ and $m$ are concatenated with respect to the channel dimension followed by a convolutional block and a vanilla CNN layer to reconstruct the image $I_{\textrm{rec}}$ that is visually identical to $I$.
3.4 Watermark Recovery
Manipulation Pools. In real-life scenarios, images inevitably suffer degradation by common post-processing operations and noises such as compression and blurring upon uploading and spreading on social media. Therefore, to satisfy the objectives, as shown in Fig. 1, we construct two manipulation pools, $P_{\textrm{common}}$ and $P_{\textrm{swap}}$, to supervise the proposed framework. On the one hand, the common manipulation pool $P_{\textrm{common}}$ includes the benign post-processing operations that are commonly seen in the real world. On the other hand, we enclose Deepfake face swapping models in the face swapping pool $P_{\textrm{swap}}$ to preserve watermark robustness when facing malicious attacks.
Decoder. We feed the post-manipulation image to the decoder for watermark recovery. The decoder follows a reverse pipeline of the encoder. In particular, after extracting features back to a large number of channels, we apply $N_{\textrm{se}}$ SE Blocks, where $N_{\textrm{se}}=N_{\textrm{dec}}+3$, to gradually expand the number of channels and seek the determinant elements that hide the watermark information. Then, $N_{\textrm{dec}}$ SE Blocks without channel modification are performed after narrowing down the number of channels. In the end, a convolutional block, a flattening operation, and a linear projection are sequentially applied to recover the watermark $m_{\textrm{rec}}$.
3.5 Objective Functions
@@
Reconstruction Loss. We embed the watermarks into images and pursue the uttermost similarity between the original and watermarked images by assigning a pixel-wise $L_{2}$ constraint on the encoder,
@@
Recovery Loss. Similarly, a bit-wise $L_{2}$ constraint is applied on the decoder for watermark recovery,
@@
where $\theta_{\textrm{dec}}$ refers to the parameters of the decoder $\textrm{Dec}(\cdot)$. At the same time, the encoder is updated alongside.
Adversarial Loss. Following the convention of robust watermarking [61, 18, 28], a discriminator $D$ with parameters $\theta_{d}$ is applied to improve visual qualities of the images following
@@
The discriminator is implemented as a binary classifier with $N_{\textrm{dis}}$ convolutional blocks, and it adversarially learns to identify clean and watermarked images on each pair of $I$ and $I_{\textrm{rec}}$ during model supervision.
Table 1: Quantitative visual quality evaluation of the watermarked images on CelebA-HQ. Information includes model name, image resolution, watermark length, PSNR, and SSIM. The best performance for each resolution is in bold and the second best performance is marked with underscore.
[…]
    Statistics adopted from the published paper due to unavailable source code for reproduction.
@@
Overall Objective Function. Coefficients $\lambda_{\textrm{enc}}$, $\lambda_{\textrm{dec}}$, $\lambda_{\textrm{adv}}$, and $\lambda_{\textrm{gen}}$ are assigned to compute the weighted sum in the overall objectives $L_{\textrm{common}}$ and $L_{\textrm{swap}}$ for $P_{\textrm{common}}$ and $P_{\textrm{swap}}$ following
@@
4 Experiment
4.1 Implementation Details
@@
Manipulation Pools. $P_{\textrm{common}}$ consists of the following manipulations for both training and testing: Dropout, Resize, Jpeg Compression, Gaussian Noise, Salt and Pepper, Gaussian Blur, and Median Blur. $P_{\textrm{swap}}$ includes SimSwap [5] solely in the training phase and adopts SimSwap, InfoSwap [10], UniFace [52], and E4S [27] with state-of-the-art performance for cross-manipulation evaluation in the testing phase. Since there are various common manipulations to be considered and they only bring noises without modifying the facial contents, we first pre-trained the model using $P_{\textrm{common}}$ for sufficient iterations. Thenceforth, we slowly fine-tuned it with the help of $P_{\textrm{swap}}$ by introducing random images that provide different facial identities for swapping onto $I_{\textrm{rec}}$.
@@
Parameters. Upon watermark generation, CelebA-HQ is adopted to derive the transform rule $\phi$ with cutoff value $c=0.5$, and we set $x_{0}=0.1$, $r=3.93$, $p=5$, and $q=11$ for the constant coefficients in Eqn.(2) and Eqn.(3). We conducted watermark encoding with lengths $64$ and $128$ for images at $128\times 128$ and $256\times 256$ resolutions, respectively, and we set $N_{\textrm{enc}}$ to $3$ and $4$, $N_{\textrm{dec}}$ to $1$ and $2$, and $N_{\textrm{dis}}$ to $3$ and $4$ for the two resolutions. In this study, we trained the model in two steps. The proposed approach is firstly trained with $P_{\textrm{common}}$ by setting $\lambda_{\textrm{enc}}=1$, $\lambda_{\textrm{dec}}=10$, and $\lambda_{\textrm{adv}}=0.01$ with a learning rate of $31e-$ 3 to guarantee the promising watermark recovery performance regarding common manipulations. Then, the trained model is further tuned on SimSwap by setting $\lambda_{\textrm{enc}}=10$, $\lambda_{\textrm{dec}}=1$, $\lambda_{\textrm{adv}}=0.01$, and $\lambda_{\textrm{gen}}=5$ with a lower learning rate of $51e-$ 5. This process mildly adjusts the framework to stay robust when facing both common and face swapping manipulations. Adam optimizers [21] are adopted through the entire training pipeline for each model component. Our model is implemented using PyTorch on $4$ Tesla V100 GPUs with a batch size of $32$.
Table 2: Quantitative comparison on CelebA-HQ at 128 and 256 resolutions regarding the bit-wise recovery accuracy of the watermarks under common manipulations and malicious face swapping manipulations. The best performance for each resolution is in bold and the second best performance is marked with underscore.
Manipulations HiDDeN  [61] $128\times 128$ MBRS  [18] $128\times 128$ RDA  [58] $128\times 128$ CIN  [28] $128\times 128$ ARWGAN  [16] $128\times 128$ Ours $128\times 128$ MBRS  [18] $128\times 128$ FaceSigns  [31] $128\times 128$ Ours $256\times 256$
Dropout 82.44% 100.00% 94.76% 100.00% 97.25% 99.99% 99.19% 55.36% 99.99%
Resize 82.01% 100.00% 99.94% 99.17% 93.73% 100.00% 100.00% 98.63% 99.08%
Jpeg 67.84% 99.49% 66.85% 96.86% 57.42% 99.48% 99.69% 82.64% 99.99%
GaussianNoise 51.36% 99.60% 60.18% 86.00% 53.45% 98.63% 58.31% 53.61% 99.97%
SaltPepper 52.30% 99.37% 66.62% 93.95% 64.55% 95.54% 65.33% 73.95% 97.54%
GaussianBlur 73.04% 100.00% 99.95% 100.00% 85.22% 100.00% 72.05% 98.68% 97.93%
MedBlur 82.72% 100.00% 99.98% 97.03% 96.66% 100.00% 98.06% 99.86% 99.01%
Average 70.24% 99.78% 84.04% 96.14% 78.33% 99.09% 84.66% 80.39% 99.07%
SimSwap [5] 50.02% 49.98% 50.00% 50.28% 52.06% 99.99% 50.00% 49.74% 99.98%
InfoSwap [10] 50.07% 50.82% 50.01% 50.60% 47.94% 99.31% 50.71% 50.00% 99.73%
UniFace [52] 54.98% 50.22% 71.15% 46.01% 59.30% 99.57% 49.98% 50.59% 94.86%
E4S [27] 49.19% 50.07% 63.03% 50.55% 49.81% 93.19% 50.07% 49.73% 93.33%
Average 51.07% 50.27% 58.55% 49.36% 52.28% 98.02% 50.19% 50.02% 96.98%
@@
Table 3: Quantitative visual quality evaluation of the watermarked images on LFW. Information includes model name, image resolution, watermark length, PSNR, and SSIM. The best performance for each resolution is in bold and the second best performance is marked with underscore.

### v3

@@ III-C Watermark Encoding
Watermark Features. As for the securely encryptedd binary identity perceptual watermark $m$ of length $l$, we first increase the length linearly via a fully connected layer followed by dimension expansion via shape rearrangement, and then leverage a convolutional block to elaborate the channel perspective. After that, we devise $N_{\textrm{enc}}$ consecutive diffusion blocks (denoted as Diff. Block) utilizing deconvolutions to smoothly diffuse the feature elements that represent $m$ to the same dimension as the width and height of $I$. Similarly, a series of $N_{\textrm{enc}}$ SE Blocks is arranged for further feature refinement for the preparation of watermark embedding.
Image Reconstruction. Lastly, features of image II and watermark $m$ are concatenated along the channel dimension, followed by a convolutional block and a vanilla CNN layer to upscale features and reconstruct the image $I_{\textrm{rec}}$ that is visually identical to $I$ but with $m$ invisibly embedded.
III-D Watermark Recovery
Manipulation Pools. In real-life scenarios, images inevitably suffer degradation due to common post-processing operations and noises such as compression and blurring upon uploading and spreading on social networks. Therefore, to satisfy the objectives, we construct two manipulation pools, namely $P_{\textrm{common}}$ and $P_{\textrm{swap}}$, to supervise the proposed framework. The common manipulation pool $P_{\textrm{common}}$ contains the benign post-processing operations that are commonly seen in the real world. On the other hand, we include Deepfake face swapping models in the face swapping pool $P_{\textrm{swap}}$ to preserve watermark robustness when facing malicious attacks.
Decoder. We feed the post-manipulation image to the decoder for watermark recovery. The decoder follows a reverse pipeline of the encoder. In particular, after extracting features back to a large number of channels, we apply $N_{\textrm{se}}$ SE Blocks to gradually expand the number of channels and seek determinant elements that hide key watermark information. Then, $N_{\textrm{dec}}$ SE Blocks without channel modification are performed after narrowing down the channel number. In the end, a convolutional block, a flattening operation, and a linear projection are sequentially applied to recover the watermark $m_{\textrm{rec}}$, which is faithfully identical to the original watermark $m$.
III-E Objective Functions
@@
Reconstruction Loss. We embed the watermarks into images and pursue the uttermost similarity between the original and watermarked images by assigning a pixel-wise $L_{2}$ constraint on the encoder following
@@
Recovery Loss. Similarly, a bit-wise $L_{2}$ constraint is applied on the decoder for watermark recovery following
@@
where $\theta_{\textrm{dec}}$ refers to the parameters of the decoder $\textrm{Dec}(\cdot)$. The encoder is updated alongside.
Adversarial Loss. Following the convention of robust watermarking [34, 35, 36], the discriminator $D$ with parameters $\theta_{d}$ is applied to improve the visual qualities of the images following
@@
The discriminator is implemented as a binary classifier with $N_{\textrm{dis}}$ convolutional blocks, and it adversarially learns to identify clean and watermarked images for each pair of II and $I_{\textrm{rec}}$ during model supervision.
@@
Overall Objective Function. Coefficients $\lambda_{\textrm{enc}}$, $\lambda_{\textrm{dec}}$, $\lambda_{\textrm{adv}}$, and $\lambda_{\textrm{gen}}$ are assigned to compute the weighted sum in the overall objectives $L_{\textrm{common}}$ and $L_{\textrm{swap}}$ for training with $P_{\textrm{common}}$ and $P_{\textrm{swap}}$ following
@@
IV Experiment
IV-A Implementation Details
@@
Manipulation Pools. $P_{\textrm{common}}$ consists of the following manipulations for both training and testing: Dropout, Resize, Jpeg Compression, Gaussian Noise, Salt and Pepper, Gaussian Blur, and Median Blur. Contrarily, $P_{\textrm{swap}}$ includes SimSwap [60] solely in the training phase and adopts SimSwap, InfoSwap [61], UniFace [62], E4S [63], DiffSwap [64], FuseAnyPart [65], Face-Adapter [66], and REFace [67] with state-of-the-art performance for cross-manipulation evaluation in the testing phase. Since there are various common manipulations to be considered and they only bring noises without modifying the facial contents, we first pre-trained the watermarking model using $P_{\textrm{common}}$ for sufficient iterations. Henceforth, we slowly fine-tuned it with the help of $P_{\textrm{swap}}$ by introducing random images that provide different facial identities to be swapped to $I_{\textrm{rec}}$.
@@
Parameters. Upon watermark generation, CelebA-HQ is adopted to derive the transform rule $\phi$ with cutoff value $c=0.5$, and we set $x_{0}=0.1$, $r=3.93$, $p=5$, and $q=11$ for the constant coefficients in Eqn.(2) and Eqn.(3). We conducted watermark encoding with lengths $64$ and $128$ for images at $128\times 128$ and $256\times 256$ resolutions, respectively, and we set $N_{\textrm{enc}}$ to $3$ and $4$, $N_{\textrm{dec}}$ to $1$ and $2$, $N_{\textrm{se}}$ to $4$ and $5$, and $N_{\textrm{dis}}$ to $3$ and $4$ for the two resolutions. In this study, we trained the model in two steps. The proposed approach is firstly trained with $P_{\textrm{common}}$ by setting $\lambda_{\textrm{enc}}=1$, $\lambda_{\textrm{dec}}=10$, and $\lambda_{\textrm{adv}}=0.01$ with a learning rate of 1e31e-3 to guarantee the promising watermark recovery performance regarding common manipulations. Then, the trained model is further tuned on SimSwap [60] by setting $\lambda_{\textrm{enc}}=10$, $\lambda_{\textrm{dec}}=1$, $\lambda_{\textrm{adv}}=0.01$, and $\lambda_{\textrm{gen}}=5$ with a lower learning rate of 1e51e-5. This process mildly adjusts the framework to stay robust when facing both common and face swapping manipulations. Adam optimizers [68] are adopted through the entire training pipeline for each model component. Our model is implemented using PyTorch on $4$ Tesla V100 GPUs with a batch size of $32$.
IV-B Experiment on CelebA-HQ
In this section, experiments are first conducted on the testing set of CelebA-HQ. Since no existing work has performed detection and source tracing on face swapping images using a single watermark to the best of our knowledge, we selected the robust watermarking frameworks (HiDDeN [34], MBRS [35], and CIN [36]), proactive robust Deepfake watermarking frameworks (RDA [40], ARWGAN [39], SepMark [^1] SepMark contains a robust and a semi-fragile watermark. Since the watermark robustness is evaluated, we picked the robust watermark of SepMark for comparison in experiments.  [21], WEvade [69], and InvisMark [70]), proactive semi-fragile Deepfake watermarking frameworks (IMD [46], MaLP [47], and FaceSigns [15]), and proactive Deepfake distortion frameworks (AntiForgery [12] and CMUA [11]) for comparison. For fairness, the publicly available trained model weights that correspond to optimal performance are directly adopted.
Figure 3: Visual effects of Deepfake face swapping image manipulations on clean and watermarked images. The first column refers to the source images that provide facial identity for face swapping. Columns 2 and 3 present the raw and watermarked target images. Every two columns of the remaining exhibit the face swapping visualizations on the raw and watermarked target images.
@@ IV-B Experiment on CelebA-HQ
TABLE I: Quantitative visual quality evaluation of the watermarked images on CelebA-HQ. Information includes model name, image resolution, watermark length, PSNR, and SSIM. The best performance for each resolution is in bold and the second best performance is marked with underscore.
[…]
HiDDeN [34] $128\times 128$ 30 33.26 0.888
MBRS [35] $128\times 128$ 30 33.01 0.775
RDA [40] $128\times 128$ 100 43.93 0.975
CIN [36] $128\times 128$ 30 43.37 0.967
IMD [46] $128\times 128$ 34.84 0.841
MaLP [47] $128\times 128$ 34.87 0.844
ARWGAN [39] $128\times 128$ 30 39.58 0.919
SepMark [21] $128\times 128$ 30 38.51 0.959
WEvade [69] $128\times 128$ 30 34.56 0.800
Ours $128\times 128$ 64 47.39 0.993
AntiForgery [12] $256\times 256$ 35.62 0.953
CMUA [11] $256\times 256$ 38.64 0.857
MBRS [35] $256\times 256$ 256 44.14 0.969
FaceSigns [15] $256\times 256$ 128 36.99 0.889
SepMark [21] $256\times 256$ 128 38.56 0.933
InvisMark [70] $256\times 256$ 100 44.20 0.997
Ours $256\times 256$ 128 45.38 0.994
TABLE II: Quantitative comparison on CelebA-HQ at the $128\times 128$ resolution regarding the bit-wise recovery accuracy of the watermarks under common manipulations and malicious face swapping manipulations. The best performance for each resolution is in bold and the second best performance is marked with underscore.
Manipulations HiDDeN [34] MBRS [35] RDA [40] CIN [36] IMD [46] MaLP [47] ARWGAN [39] SepMark [21] WEvade [69] Ours
Dropout 82.44% 99.99% 94.76% 99.99% 0.5573 0.4532 97.25% 99.99% 78.02% 99.99%
Resize 82.01% 99.99% 99.94% 99.17% 0.4092 0.4562 93.73% 99.76% 69.67% 99.99%
Jpeg 67.84% 99.49% 66.85% 96.86% 0.5084 0.1379 57.42% 98.14% 66.69% 99.48%
GaussianNoise 51.36% 99.60% 60.18% 86.00% 0.4497 0.4458 53.45% 59.31% 64.87% 98.63%
SaltPepper 52.30% 99.37% 66.62% 93.95% 0.4422 0.2804 64.55% 99.90% 55.79% 95.54%
GaussianBlur 73.04% 99.99% 99.95% 99.99% 0.3609 0.4581 85.22% 99.30% 65.58% 99.99%
MedBlur 82.72% 99.99% 99.98% 97.03% 0.4120 0.4544 96.66% 99.98% 68.12% 99.99%
Average 70.24% 99.78% 84.04% 96.14% 0.4485 0.3837 78.33% 93.77% 66.96% 99.09%
SimSwap [60] 50.02% 49.98% 50.00% 50.28% 0.0035 0.7228 52.06% 59.02% 59.25% 99.99%
InfoSwap [61] 50.07% 50.82% 50.01% 50.60% 0.5261 0.1489 47.94% 76.62% 66.51% 99.31%
UniFace [62] 54.98% 50.22% 71.15% 46.01% 0.4093 0.4467 59.30% 66.13% 51.50% 99.57%
E4S [63] 49.19% 50.07% 63.03% 50.55% 0.0071 0.6408 49.81% 79.72% 74.06% 93.19%
DiffSwap [64] 40.53% 43.48% 61.94% 47.15% 0.5298 0.4935 58.30% 72.73% 78.77% 80.10%
FuseAnyPart [65] 73.57% 88.95% 48.25% 66.88% 0.4891 0.6010 69.58% 90.73% 80.56% 91.00%
Face-Adapter [66] 73.12% 76.74% 50.42% 68.57% 0.3799 0.6109 71.84% 79.10% 75.67% 82.26%
REFace [67] 48.17% 89.03% 49.86% 88.56% 0.6206 0.6240 80.01% 90.52% 76.87% 94.61%
Average 54.96% 62.41% 55.58% 58.58% 0.3707 0.5361 61.11% 76.83% 70.40% 92.50%
  • $\dagger$ IMD and MaLP are specially evaluated via cosine similarity according to the original paper. Higher values denote better watermark robustness.
TABLE III: Quantitative comparison on CelebA-HQ at the $256\times 256$ resolution regarding the bit-wise recovery accuracy of the watermarks under common manipulations and malicious face swapping manipulations. The best performance for each resolution is in bold and the second best performance is marked with underscore.
Manipulations MBRS [35] FaceSigns [15] SepMark [21] InvisMark [70] Ours
Dropout 99.19% 55.36% 99.99% 99.59% 99.99%
Resize 99.99% 98.63% 99.99% 99.99% 99.08%
Jpeg 99.69% 82.64% 99.94% 99.91% 99.99%
GaussianNoise 58.31% 53.61% 52.71% 99.79% 99.97%
SaltPepper 65.33% 73.95% 99.99% 87.27% 97.54%
GaussianBlur 72.05% 98.68% 99.92% 99.99% 97.93%
MedBlur 98.06% 99.86% 99.99% 99.99% 99.01%
Average 84.66% 80.39% 93.22% 98.13% 99.07%
SimSwap [60] 50.00% 49.74% 54.07% 52.92% 99.98%
InfoSwap [61] 50.71% 50.00% 83.09% 61.70% 99.73%
UniFace [62] 49.98% 50.59% 56.34% 53.06% 94.86%
E4S [63] 50.07% 49.73% 77.61% 60.79% 93.33%
DiffSwap [64] 77.28% 52.06% 76.95% 52.87% 93.43%
FuseAnyPart [65] 84.33% 49.99% 98.49% 52.99% 98.53%
Face-Adapter [66] 74.10% 50.59% 82.54% 53.17% 87.23%
REFace [67] 75.84% 70.68% 78.33% 53.13% 94.83%
Average 64.04% 52.92% 75.93% 55.08% 95.24%
IV-B1 Qualitative Evaluation
We visualized randomly drafted images to validate the visual quality of the watermarked images. As exhibited in Fig. 2, despite knowing the existence of the watermarks, differences between the raw and watermarked images in the first and second columns, respectively, are visually imperceptible. Starting from column 3, we illustrated the visual effects of the common manipulations on the watermarked images, one for each column. In Fig. 3, the face swapping effects for the GAN-based [^2] Diffusion-based face swapping results are omitted since they contain randomness, which naturally leads to slightly different outputs from two identical inferences even without introducing watermarks. generative models are exhibited in every two columns starting from column 4. The source images that provide the desired synthetic facial identities for each row are placed in the first column. Columns 2 and 3 display the raw and watermarked target images, while every pair of the remaining columns illustrates the face swapping results on the raw and watermarked target images, respectively. It is worth noting that no distortion can be observed in the face swapping results using watermarked target images. In other words, the encoded watermarks do not affect the visual qualities of the synthetic faces, preserving the potential for benign usages of Deepfake face swapping in the industry.
IV-B2 Quantitative Evaluation
Visual Quality. We employed the average peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) [71] between the raw and watermarked images to evaluate the visual quality of watermarked images. As Table I reports, while the goal of embedding watermarks into raw images imperceivably stands for all contrastive methods, our proposed approach achieves state-of-the-art visual quality in terms of PSNR and SSIM at both resolution levels, except for ranking a competitive second place for SSIM at the $256\times 256$ resolution. Meanwhile, for the $128\times 128$ resolution, RDA [40] and CIN [36] appear to have reasonable PSNR values above $43$ and SSIM values above $0.95$. Although slightly behind in terms of PSNR, ARWGAN [39] and SepMark [21] perform well with SSIM values above $0.9$. For the $256\times 256$ resolution, MBRS [35] and InvisMark [70] are promising for both PSNR and SSIM metrics, where InvisMark exhibits competitive performance compared to our approach. On the other hand, it can be observed that, although pursuing the same objective in visual quality, the proactive distortion models, AntiForgery [12] and CMUA [11], have revealed relatively unsatisfactory performance, since noises and perturbations are directly added to raw images, while watermarking approaches execute image reconstructions.
where $d_{\textrm{ham}}$ computes the hamming distance [72] between the recovered and original watermarks $m_{\textrm{rec}}$ and $m$, respectively, and $l$ refers to the watermark length. In the experiment, each type of manipulation in $P_{\textrm{common}}$ and $P_{\textrm{swap}}$ is performed on the watermarked image $I_{\textrm{rec}}$ before extracting the watermark via the decoder.
The watermark recovery performance in the comparative experiment is reported in Table II and Table III. Specifically, considering the common manipulations, our method outperforms HiDDeN [34], RDA [40], CIN [36], ARWGAN [39], SepMark [21], and WEvade [69] at the $128\times 128$ resolution with a better average accuracy, and our performance is competitive against MBRS [35] with an acceptable average accuracy above $99\%$, achieving the second best statistic. However, it can be observed from Table I that MBRS heavily suffers from visual quality issues with unsatisfactory PSNR and SSIM values ($33.01$ and $0.775$) in order to achieve the best bit-wise watermark recovery accuracy in Table II. As for the $256\times 256$ resolution, in Table III, we reached an average accuracy of $99.07\%$ on the common manipulations, outperforming existing state-of-the-art methods. On the contrary, although accomplishes reasonable visual quality as listed in Table I, MBRS experiences a large performance damping on the common manipulations in Table III with respect to the watermark recovery accuracy compared to its $128\times 128$ resolution version. Meanwhile, although having an average robustness slightly lower than ours, InvisMark [70] achieves a promising second-best value of $98.13\%$. On the other hand, IMD [46], MaLP [47], and FaceSigns [15] are semi-fragile approaches, where the former two justify watermark recovery performance by the cosine similarity between recovered and original watermarks. Unfortunately, they each fail to stay robust when facing some common manipulations, leading to relatively far gaps from the perfect cosine similarity value ($1.0$) and accuracy ($100.00\%$). More findings are discussed when evaluating them against face swapping manipulations, as they are supposed to be fragile.
Watermark Recovery on Face Swapping. In this paper, the face swapping manipulation pool contains merely SimSwap [60] in the training phase, and we validated the watermark recovery performance for both in- and cross-manipulation settings by introducing GAN-based face swapping algorithms SimSwap [60], InfoSwap [61], UniFace [62], and E4S [63], and diffusion-based face swapping algorithms DiffSwap [64], FuseAnyPart [65], Face-Adapter [66], and REFace [67] in the testing phase. As shown in Table II and Table III, our model consistently recovers the desired watermark messages at average accuracies of $92.50\%$ and $95.24\%$ for $128\times 128$ and $256\times 256$ resolutions, respectively, while HiDDeN, MBRS, RDA, CIN, ARWGAN, and InvisMark obtain accuracies around $60\%$ against most synthetic models. WEvade performs better than the former, but still struggles at around $70\%$ on average. Moreover, it can be concluded from the cross-manipulation tests on unseen synthetic models that, by simply tuning with SimSwap, the hidden watermarks are pleasantly maintained by our approach since the goals for different face swapping algorithms are ultimately identical. The semi-fragile watermarking framework, FaceSigns, obtains low watermark recovery accuracies as designed when facing synthetic manipulations that are expected to destroy the embedded watermarks. However, its performance regarding the common manipulations is unexpectedly unsatisfactory besides illustrating low visual quality as reported in Table I, leading to an average recovery accuracy of $80.39\%$. Meanwhile, IMD and MaLP exhibit rare robustness when facing one or two particular face swapping algorithms, while having even higher cosine similarities (e.g., $0.7228$ against SimSwap for MaLP) than those when facing common manipulations. Lastly, although preserves outstanding robustness against most common manipulations by obtaining $93.77\%$ and $93.22\%$ average accuracies, SepMark fails to recover the original robust watermarks with only $76.83\%$ and $75.93\%$ average accuracies for the two resolutions, respectively.
Discussion. In general, Salt and Pepper and Gaussian Noise are the most challenging common manipulations that cause fluctuations in the watermark recovery accuracies for most watermarking algorithms as summarized in Table II and Table III. This is consistent with the visualization results in Fig. 2 (col. 6 & 9) such that these two manipulations add a plethora of perceivable noises to the images and thus heavily jeopardize the hidden watermarks. On the other hand, the contrastive robust watermarking frameworks fail to consistently maintain robustness when facing state-of-the-art face swapping models. As a result, our proposed model is the only one that maintains the bit-wise watermark recovery accuracies above 95% for all common manipulations and above $80\%$ and $85\%$ at $128\times 128$ and $256\times 256$ resolutions, respectively, for all face swapping manipulations, while the contrastive methods are heavily affected by at least one of the manipulations.
Moreover, although successfully maintains reasonable robustness in general, our method is observed to be affected more when facing the diffusion-based algorithms, with as low as $80.10\%$ regarding robustness against DiffSwap [64]. This performance damping is caused due to the way that diffusion models generate outputs. In particular, DiffSwap edits the target image in a fixed number of iterations, where each iteration derives an intermediate result, leading to possible information loss regarding the embedded watermark. Consequently, the ultimate face swapping results suffer certain levels of damage in the underlying watermarks, which explains the statistical results in Table II and Table III. Another observation appears to be that the watermarking algorithms generally perform better at the higher resolution against diffusion-based face swapping models. This is possibly because diffusion-based models are mostly designed to deal with higher resolutions. Meanwhile, some diffusion-based models edit specific facial parts rather than whole face reconstruction, which then leads to easier conditions for watermarks to survive. For instance, FuseAnyPart [65] crops and substitutes only the eyes, nose, and mouth areas during the face swapping process, allowing all watermarking algorithms to perform better than when they defend against other diffusion-based models. As a result, despite oscillations in the accuracies against different face swapping models, satisfactory watermark robustness can still be observed for our approach compared to the state-of-the-art ones.
TABLE IV: Quantitative visual quality evaluation of the watermarked images on LFW. Information includes model name, image resolution, watermark length, PSNR, and SSIM. The best performance for each resolution is in bold and the second best performance is marked with underscore.